# T-3DGS: Removing Transient Objects for 3D Scene Reconstruction

### v1

@@ Abstract
We propose a novel framework to remove transient objects from input videos for 3D scene reconstruction using Gaussian Splatting. Our framework consists of the following steps. In the first step, we propose an unsupervised training strategy for a classification network to distinguish between transient objects and static scene parts based on their different training behavior inside the 3D Gaussian Splatting reconstruction. In the second step, we improve the boundary quality and stability of the detected transients by combining our results from the first step with an off-the-shelf segmentation method. We also propose a simple and effective strategy to track objects in the input video forward and backward in time. Our results show an improvement over the current state of the art in existing sparsely captured datasets and significant improvements in a newly proposed densely captured (video) dataset. More results and code are available at https://transient-3dgs.github.io.
Figure 1: Recent state-of-the-art methods, e.g. SpotLessSplats [34], often fail to correctly identify masks for transient / semi-transient objects. Our proposed T-3DGS method successfully locates all transient distractors, extracts clean masks, and propagates them between frames. By masking transient objects, T-3DGS can produce clean reconstruction results with high-fidelity novel view synthesis on a given sequence of images or videos captured in the wild.
@@ 1 Introduction
Novel view synthesis and 3D scene reconstruction from multiple 2D images or videos have emerged as rapidly evolving research areas within computer vision. Neural Radiance Fields (NeRF) [21] and, more recently, 3D Gaussian Splatting (3DGS) [12] have shown remarkable improvements in novel view synthesis on complex scenes. NeRF implicitly represents the scene as a volumetric function, and 3DGS explicitly represents it as a set of 3D Gaussians. Both approaches produce high-quality realistic images. There are multiple follow-up works for diverse downstream applications including 3D scene reconstruction [44, 16, 9], 3D synthesis [25, 48, 39], semantic and language integration into 3D representations [37, 13, 36], etc.
@@
Both 3D Gaussian Splatting and NeRF optimize 3D scene reconstruction using photometric losses. High-quality results are achieved under the assumption that the captured scene is completely static and does not include any distractors, such as moving objects (i.e. transient objects), shadows, lightning changes, etc. However, in real-world scenarios, such an assumption can hardly be satisfied. The recordings, even when captured carefully, often contain moving people, cars, or other dynamic objects with their shadows, especially in locations that are tourist landmarks. Ignoring distractors during scene optimization leads to undesired blurring effects and floating artifacts. At the same time, removing such distractors from the captured recordings is challenging and limits NeRF and 3DGS widespread usage. The process of manual annotations of distractors is labor-intensive. Another approach is to utilize pre-trained segmentation models to locate transient distractors. This approach has two main limitations: 1) it needs prior knowledge of transients as a semantic class, and 2), more importantly, existing segmentation models cannot distinguish between static and dynamic objects of the same semantic class. Additionally, we would like to identify semi-transient objects in recordings and remove them from the scene. We define a semi-transient object as an object that has both dynamic and static states during the capturing process, e.g. a pushed chair stops after some time and becomes a fully static object. Therefore, we need more robust identification methods for transient and semi-transient distractors throughout the captured recordings.
In our work, we introduce a novel T-3DGS approach to 3D static scene reconstruction with 3DGS from a monocular video taken in uncontrolled setting. Our method comprises two main components: an unsupervised transient detector and a transient mask propagation framework. Firstly, as it was observed previously [31, 34], relying only on image residuals in transient identification is not a robust approach, e.g., due to appearance changes and color similarity between transient objects and background. We need more robust features, which are semantically consistent as well as novel architecture and loss design. Our proposed transient network is trained concurrently on both rendered (reconstructed) images and reference images through novel robust regularization losses. It utilizes semantically consistent features and is capable of producing more accurate masks with fewer misidentifications of transient objects, and significantly reduces the number of falsely identified objects as dynamic objects. Secondly, we have shown in our experiments that concurrent works [34, 43] fail to identify the semi-transient distractors and remove them from the scene as shown in Fig. 1. We further introduce our mask propagation framework that allows the extraction of object-aware masks that are consistent throughout the video and are especially helpful in improving the consistency of semi-transient distractor masks. Overall, our method remains robust for any type of distractors present in the scene. Lastly, we have collected a novel T-3DGS dataset containing several challenging scenes filled with semi-transient and slow-moving objects with dense frame rates. We evaluate our approach on casually captured scenes [31, 33] and on our semi-transient dataset. Our method consistently outperforms state-of-the-art approaches in reconstruction quality.
Our key contributions include:
  • A novel transient network to identify dynamic distractors utilizing features from a visual foundation model.
@@
  • A novel video object segmentation module that robustly tracks objects throughout the video regardless of the frame rate and semi-transient behavior.
@@
  • A novel challenging dataset that contains diverse scenes with semi-transient distractors and scenes with slow-moving objects for evaluation.
@@
  • Our proposed method presents new state-of-the-art performance on robust reconstruction of the static scene parts on standard benchmark datasets.
@@ 2 Related Work
We briefly review Neural Radiance Fields and 3D Gaussian Splatting works with a focus on removing non-static distractors in the scene.
@@
Neural Radiance Fields (NeRFs) [21] have emerged as a widely adopted method for high-quality scene reconstruction and novel view synthesis of 3D scenes. There are several follow-up extensions of NeRFs for fast training and inference [38, 22, 3], novel view synthesis on dynamic scenes [26, 42, 24], training on single or limited views [11, 30]. The seminal paper 3D Gaussian Splatting [12] employs Gaussian primitives to model scenes instead of relying on continuous volumetric representations. This method has recently gained popularity as a faster alternative to NeRFs. Follow-up works have extended 3DGS in many directions, for instance in dynamic [47, 52] and static [20, 50] scenes, utilizing additional information such as depth [55].
@@
Handling Distractors in NeRFs. NeRF-W [19] and RobustNeRF [33] are two pioneering works similarly approaching the problem of handling transient distractors by utilizing photometric error. NeRF-W reconstructs both static background and transients combined with a data-dependent uncertainty field. RobustNeRF utilizes Iteratively Reweighted Least Squares for transient object identification and removal. Both methods rely directly on color residual supervision and often misclassify transient objects and backgrounds with similar colors. Additionally, they both require careful hyper-parameters tuning. Another alternative approach is NeRFBusters [46], which uses a data-driven prior to remove artifacts. NeRF On-the-go [31] utilizes DINOv2 features [23] to identify and eliminate distractors by predicting uncertainties through a shallow MLP and can deal with more complex scenes than RobustNeRF. NeRF-HuGS [2] utilizes two types of heuristics: 1) COLMAP-based [35] features combined with SAM [14] and 2) residual-based heuristics to identify and remove transient distractors. Their method is not robust to heavy transient distractions because both heuristics are not stable under such conditions, as shown in [31].
Extracting Features from Vision Foundation Models. Vision Foundation Models (VFMs) are trained on large-scale data to perform strong generalization ability to unseen domains or novel tasks. For example, there are a set of task-specific foundation models for image classification  [7, 41, 17], for segmentation [14, 29], and depth prediction [51] tasks. Task-agnostic models trained through self-distillation like DINO [1, 23] learn features that can be generalized for multiple vision tasks. CLIP [27] is a weakly-supervised multi-modal model that learns visual concepts from natural language supervision by contrastive pre-training on image-text pairs and performs zero-shot transfer to a wide range of downstream tasks. The features from Diffusion Models [10, 32] have also been used for semantic or keypoint correspondence [40, 54, 18]. In our work, we aim to utilize features from foundation models to identify transient distractors more robustly. We propose to use DINOv2 features [23] because they are semantically meaningful and spatially consistent, which is crucial for reliable distractor identification.
@@
Video Object Segmentation. The goal of semi-supervised VOS is to identify when an object appears for the first time and then track it throughout the video. There are several recent approaches based on transformers [4, 5]. However, current methods suffer from mask inconsistencies throughout the video, especially when objects disappear from the video and reappear later. Additionally, current methods assume that the input is provided with a high frame rate, and these methods are not stable when the frame rate is low. In our work, we address these shortcomings.
@@
Handling Distractors in 3DGS. There are multiple works addressing training 3DGS on unconstrained in-the-wild photo collections. SWAG [6] improves robustness of 3DGS by learning an appearance embedding space and image-dependent opacity variations to handle transient objects better. Gaussians in the Wild (GS-W) [53] utilizes CNN features to capture dynamic and intrinsic appearances from a reference image. Wild-GS [49] explicitly learns appearance embeddings by sampling the triplane from the reference image. Robust 3DGS [43] proposes a self-supervised approach to identify transient distractors by utilizing image residuals and leveraging a pre-trained segmentation network to produce object-aware masks. SpotLessSplats [34] proposes a method to identify transient objects by utilizing pre-computed feature maps from a foundation model [40] coupled with a robust optmization of 3DGS. These works [6, 53, 49, 43, 34] suffer from 1) hyper-parameter tuning, i.e., threshold parameters, 2) predicting inaccurate transient masks across the video and 3) relying on image residuals that result in falsely detecting transients as shown in Fig. 1. In our approach we try to address the limitations of the current works by identifying transients more accurately and consistently between video frames.
@@ 3 Method
Figure 2: The architecture of the proposed method. We use a modified version of 3D Gaussian Splatting (in green), with masked loss term $\mathcal{L}_{\text{masked}}$ specified in (Eq. 6). Every iteration, we begin by rendering a reconstruction of a randomly sampled reference image. Then we calculate RGB color residuals (i.e $L_{1}$ difference of those images), and DINOv2 features of both images. These features are then fed to our TMP model (in purple) that predicts binary masks for both reference and rendered images (shown in green on the actual images for clarity). We then calculate the loss specified in (Eq. 1) and update the parameters of the TMP via backpropagation, denoted with dashed lines. Notably, we detach rendered image masks to prevent the TMP from learning on features of the rendered image. Additionally, we use a SAM-based mask propagation module, as it further increases the consistency and sharpness of masks.
We propose a novel approach for reconstructing static scenes from unconstrained videos containing dynamic objects using 3D Gaussian Splatting (3DGS). Our method, illustrated in Fig. 2, introduces two key components to handle dynamic objects effectively: (1) transient mask predictor (TMP) and (2) transient mask refiner (TMR). The transient area detection component, implemented through our transient mask learning predictor, identifies regions containing dynamic objects by predicting per-pixel probabilities using semantic features, enhanced with a novel consistency regularization term. This regularization reduces false positive predictions while maintaining accurate detection of dynamic objects. The transient mask propagator component improves transient detection in both the spatial and temporal domain by leveraging SAM2 [29] to propagate transient masks across different frames, enabling artifact-free reconstruction.
@@
We base our method on 3DGS [12]. Given a set of posed images $\{I_{n}\}_{n=1}^{N},I_{n}\in\mathbb{R}^{H\times W\times C}$, 3DGS represents a 3D scene as a set of anisotropic Gaussians $\{\mathcal{G}_{i}\}$, where each Gaussian is represented by its position (mean) $\mu_{i}$, a positive semi-definite covariance matrix $\Sigma_{i}$, an opacity $\alpha_{i}$, and a view-dependent appearance component (color) parametrized by spherical harmonics (SH) [28]. 3DGS representation is learned through optimization of Gaussian parameters via stochastic gradient descent.
@@
Each 3D Gaussian is projected onto the image plane through a differentiable rasterization process to render an image from a specific viewpoint. First, the 3D Gaussian’s covariance matrix $\Sigma_{i}$ is projected to obtain a 2D covariance matrix $\Sigma^{\prime}_{i}$ in screen space: $\Sigma^{\prime}_{i}=JW\Sigma_{i}W^{T}J^{T}$, where $W$ is the perspective transformation matrix and $J$ is the Jacobian of the projection matrix. The projected 2D Gaussian’s contribution to each pixel $(x,y)$ is then computed as: $\alpha_{i}=\exp(-\frac{1}{2}(p-\mu^{\prime}_{i})^{T}(\Sigma^{\prime}_{i})^{-1}(p-\mu^{\prime}_{i}))$, where $p$ is the pixel coordinate and $\mu^{\prime}_{i}$ is the projected mean. The final color at each pixel is obtained through alpha compositing of all contributing Gaussians, sorted by depth: $C=\sum_{i=1}^{M}T_{i}\alpha_{i}c_{i}$, where $T_{i}=\prod_{j<i}(1-\alpha_{j})$ is the accumulated transmittance, $c_{i}$ is the view-dependent color computed from spherical harmonics coefficients, and $M$ is the number of Gaussians contributing to the pixel.
3.2 Transient Mask Prediction
@@
Given the input images $\{I_{n}\}_{n=1}^{N}$, the goal is to optimize unsupervised TMP through 3DGS reconstruction to identify transient distractors without explicit supervision as shown in Fig. 2. The training process consists of two steps: (1) optimizing the 3DGS parameters while keeping the TMP frozen, and (2) optimizing the TMP while keeping the 3DGS parameters frozen. After each iteration, both models’ weights are updated. The TMP is trained to identify transient objects without explicit supervision, purely from the reconstruction objectives. Several recent works [6, 34, 15, 8, 49] follow a similar approach, which has demonstrated the effectiveness of such optimization for handling dynamic scenes.
@@
Following [15, 34], we reformulate the transient detection problem as a semantic feature classification task rather than a direct pixel-level prediction. This approach leverages pre-trained foundation models to extract rich semantic features from images, enabling our system to make decisions based on high-level semantic understanding rather than solely relying on color information. This semantics-aware approach proves to be more robust in distinguishing between static and transient objects compared to traditional color-based methods. One of our main contributions is the unsupervised consistency loss in Eq. 3.

### v2

@@ Abstract
Transient objects in video sequences can significantly degrade the quality of 3D scene reconstructions. To address this challenge, we propose T-3DGS, a novel framework that robustly filters out transient distractors during 3D reconstruction using Gaussian Splatting. Our framework consists of two steps. First, we employ an unsupervised classification network that distinguishes transient objects from static scene elements by leveraging their distinct training dynamics within the reconstruction process. Second, we refine these initial detections by integrating an off-the-shelf segmentation method with a bidirectional tracking module, which together enhance boundary accuracy and temporal coherence. Evaluations on both sparsely and densely captured video datasets demonstrate that T-3DGS significantly outperforms state-of-the-art approaches, enabling high-fidelity 3D reconstructions in challenging, real-world scenarios. More results and code are available at https://transient-3dgs.github.io/
Figure 1: Existing state-of-the-art methods, such as SpotLessSplats [22], often struggle to correctly identify transient and semi-transient objects, leading to artifacts in 3D scene reconstruction. Our proposed T-3DGS method accurately detects all transient distractors, generates clean masks, and propagates them across frames. By effectively masking transient objects, T-3DGS enables high-fidelity novel view synthesis and significantly improves reconstruction quality from real-world image and video sequences.
22footnotetext: Correspondence to: a.komarichev@skoltech.ru.
@@ 1 Introduction
Novel view synthesis and 3D scene reconstruction from multiple 2D images or videos are critical, rapidly evolving areas in computer vision. Neural Radiance Fields (NeRF) [15] and 3D Gaussian Splatting (3DGS) [9] have shown remarkable improvements in novel view synthesis on complex scenes. NeRF implicitly represents the scene as a volumetric function, and 3DGS explicitly represents it as a set of 3D Gaussians. Both approaches produce high-quality realistic images. There are multiple follow-up works for diverse downstream applications, including 3D scene reconstruction [29, 13, 8], 3D synthesis [17, 31, 26], semantic and language integration into 3D representations [25, 10, 24].
@@
Both 3D Gaussian Splatting and NeRF optimize 3D scene reconstruction using photometric losses. High-quality results are achieved under the assumption that the captured scene is completely static and does not include any distractors, such as moving objects (i.e. transient objects), shadows, lightning changes, etc. In real-world scenarios, this assumption can hardly be satisfied. Even when carefully captured, recordings often contain moving people, cars, or other dynamic objects with their shadows, especially in locations that are tourist landmarks. Ignoring distractors during scene optimization results in undesired blurring effects and floating artifacts. At the same time, removing such distractors from the captured recordings is challenging and limits the widespread usage of NeRF and 3DGS. Manually annotating distractors is labor intensive. Another approach is to utilize pre-trained segmentation models to locate transient distractors. This approach has two main limitations: 1) it needs prior knowledge of transients as a semantic class, and 2), more importantly, existing segmentation models cannot distinguish between static and dynamic objects of the same semantic class. Additionally, we would like to identify semi-transient objects in recordings and remove them from the scene. We define a semi-transient object as an object that has both dynamic and static states during the capturing process, e.g. a pushed chair stops after some time and becomes a fully static object. Therefore, we need more robust identification methods for transient and semi-transient distractors throughout the captured recordings.
We introduce T-3DGS, a novel approach for 3D static scene reconstruction from monocular video in uncontrolled settings. Our method includes an unsupervised transient detector and a transient mask propagation framework. Relying solely on image residuals for transient identification is unreliable due to issues such as appearance changes and color similarity to the background [20, 22]. To address this issue, we develop a divergence-based technique on top of the uncertainty modeling approach [12] to detect transients. It helps improve mask accuracy and significantly reduce misclassifications of transient objects.
Our experiments show that concurrent works [22, 28] fail to remove semi-transient distractors (Fig. 1). We introduce a mask propagation framework for extracting object-aware masks that improve consistency in case of semi-transient distractors. Our method remains robust to all types of distractors. Additionally, we present the novel T-3DGS dataset with challenging scenes featuring semi-transient and slow-moving objects. Evaluations on both casual scenes [20, 21] and our dataset show our method outperforms state-of-the-art approaches in reconstruction quality.
Our key contributions, which together ensure consistent detection and removal of transient objects for improved 3D reconstruction, include:
  • Generalized uncertainty modeling for efficient transient object identification;
  • A divergence-based approach that leverages semantic consistency between reference and reconstructed frames for identifying transient objects;
@@
  • A robust video object segmentation module that tracks objects across varying frame rates and semi-transient behaviors;
@@
  • A challenging new dataset featuring diverse scenes with semi-transient distractors and slow-moving objects;
@@
  • State-of-the-art performance on benchmark datasets for robust static scene reconstruction.
@@ 2 Related Work
We provide a brief review of the works on Neural Radiance Fields and 3D Gaussian Splatting with a focus on removing non-static distractors in the scene.
@@
Neural Radiance Fields (NeRFs) [15] are widely adopted methods for high-quality scene reconstruction and novel view synthesis of 3D scenes. The seminal work 3D Gaussian Splatting [9] employs Gaussian primitives to model scenes instead of relying on continuous volumetric representations. This method has recently gained popularity as a faster alternative to NeRFs.
@@
Handling Distractors in NeRFs. NeRF-W [14] and RobustNeRF [21] are two pioneering works approaching the problem in a similar way. NeRF-W reconstructs both static background and transients combined with a data-dependent uncertainty field. RobustNeRF utilizes Iteratively Reweighted Least Squares for transient object identification and removal. Both methods rely on color residual supervision and frequently misclassify transient objects and backgrounds that share similar colors. Additionally, they both require careful hyper-parameters tuning. NeRF On-the-go [20] utilizes DINOv2 features [16] to identify and eliminate distractors by predicting uncertainties through a shallow MLP and can deal with more complex scenes than RobustNeRF. NeRF-HuGS [2] utilizes two types of heuristics: 1) COLMAP-based [23] features combined with SAM [11] and 2) residual-based heuristics to identify and remove transient distractors. Their method lacks robustness to heavy transient distractions, as both heuristics are unstable under such conditions, as demonstrated in [20].
Extracting Features from Vision Foundation Models. Vision Foundation Models (VFMs) are trained on large-scale data, enabling strong generalization to unseen domains or novel tasks. Task-agnostic models trained through self-distillation like DINO [1, 16] learn features that can be generalized for multiple vision tasks.
@@
Video Object Segmentation. The goal of semi-supervised VOS is to identify when an object appears for the first time and then track it throughout the video. Several recent approaches based on transformers [3, 4] have been proposed. However, current methods suffer from mask inconsistencies, particularly when objects disappear and reappear in the video. Additionally, these methods assume that the input has a high frame rate, and they become unstable when the frame rate is low. In our work, we address these shortcomings.
Figure 2: Overview of the Proposed T-3DGS Architecture. We introduce a modified version of 3D Gaussian Splatting, incorporating a masked loss term $\mathcal{L}_{\text{masked}}$ as described in Eq. 13. In each iteration, we start by rendering a reconstruction of a randomly sampled reference image. We compute residuals, along with DINOv2 features from both the ground truth and rendered images. These features are then fed to our RUP model to predict the per-pixel covariance matrix for both images. We calculate binary masks based on the divergence of these distributions (as specified in Eq. 10). Subsequently, we compute the likelihood as described in Eq. 7 and update the parameters of the RUP model via backpropagation, as indicated by the dashed lines. Additionaly, for some scenes, we incorporate a SAM-based mask refiner module (TMR), which further enhances the consistency and sharpness of the masks.
@@
Handling Distractors in 3DGS. Several works address the training of 3DGS on unconstrained, in-the-wild photo collections. SWAG [5] improves robustness of 3DGS by learning an appearance embedding space and image-dependent opacity variations to handle transient objects better. Gaussians in the Wild (GS-W) [33] utilizes CNN features to capture dynamic and intrinsic appearances from a reference image. Wild-GS [32] explicitly learns appearance embeddings by sampling the triplane from the reference image. Robust 3DGS [28] proposes a self-supervised approach to identify transient distractors by utilizing image residuals and leveraging a pre-trained segmentation network to produce object-aware masks. SpotLessSplats [22] proposes a method to identify transient objects by utilizing pre-computed feature maps from a foundation model [27] coupled with a robust optmization of 3DGS. These works [5, 33, 32, 28, 22] suffer from: 1) the need for hyper-parameter tuning, such as threshold parameters; 2) inaccurate prediction of transient masks across the video; and 3) reliance on image residuals, leading to the false detection of transients, as shown in Fig. 1. In our approach, we aim to address the limitations of the current works by identifying transients more accurately and consistently across video frames.
@@ 3 Method
We propose a novel approach to reconstructing static scenes from unconstrained videos that contain dynamic objects, utilizing 3D Gaussian Splatting (3DGS). Our method, illustrated in Fig. 2, introduces two key components designed to handle dynamic objects effectively: (1) reconstruction uncertainty predictor (RUP), and (2) transient mask refiner (TMR). The transient area detection component, implemented through our transient mask learning predictor, identifies regions containing dynamic objects by predicting per-pixel probabilities using semantic features. The transient mask refiner improves transient detections in both spatial and temporal domains by leveraging SAM2 [19] to propagate transient masks across multiple frames, facilitating artifact-free reconstruction.
@@
Our method is based on 3DGS [9]. Given a set of posed images $\{I_{n}\}_{n=1}^{N},I_{n}\in\mathbb{R}^{H\times W\times C}$, 3DGS represents a 3D scene as a set of anisotropic Gaussians $\{\mathcal{G}_{i}\}$, where each Gaussian is represented by its position (mean) $\mu_{i}$, a positive semi-definite covariance matrix $\Sigma_{i}$, an opacity $\alpha_{i}$, and a view-dependent appearance component (color) parametrized by spherical harmonics (SH) [18]. 3DGS representation is learned through optimization of Gaussian parameters via stochastic gradient descent.
@@
Each 3D Gaussian is projected onto the image plane through a differentiable rasterization process to render an image from a specific viewpoint. First, the 3D Gaussian’s covariance matrix $\Sigma_{i}$ is projected to obtain a 2D covariance matrix $\Sigma^{\prime}_{i}$ in screen space: $\Sigma^{\prime}_{i}=JW\Sigma_{i}W^{T}J^{T}$, where $W$ is the perspective transformation matrix and $J$ is the Jacobian of the projection matrix. The contribution of projected 2D Gaussian to each pixel $(x,y)$ is computed as: $\alpha_{i}=\exp(-\frac{1}{2}(p-\mu^{\prime}_{i})^{T}(\Sigma^{\prime}_{i})^{-1}(p-\mu^{\prime}_{i}))$, where $p$ is the pixel coordinate and $\mu^{\prime}_{i}$ is the projected mean. The final color at each pixel is obtained by alpha compositing the contributions from all Gaussians, sorted by depth: $C=\sum_{i=1}^{M}T_{i}\alpha_{i}c_{i}$, where $T_{i}=\prod_{j<i}(1-\alpha_{j})$ is the accumulated transmittance, $c_{i}$ is the view-dependent color computed from spherical harmonics coefficients, and $M$ is the number of Gaussians contributing to the pixel.
3.2 Reconstruction Uncertainty Prediction
@@
Given the input images $\{I_{n}\}_{n=1}^{N}$, the goal is to optimize the unsupervised reconstruction uncertainty predictor RUP through 3DGS reconstruction to identify transient distractors without explicit supervision as shown in Fig. 2. Following prior research [14, 20, 12] we employ uncertainty modeling techniques, with significant modifications. RUP is trained to identify transient objects without explicit supervision, purely from the reconstruction objectives. Several recent works [5, 22, 12, 7, 32] follow a similar approach, demonstrating the effectiveness of this optimization in handling dynamic scenes. As in WildGaussians [12] (and, in contrast to NeRF counterparts [14, 20]) every iteration we update both Gaussian Splatting and RUP weights. Additionally, we detach masks when updating Gaussian Splatting and detach reconstructed images when updating RUP.
@@
Following [12, 22], we reformulate the transient detection problem as a semantic feature classification task. This approach leverages pre-trained foundation models to extract rich semantic features from images. By doing so, it enables our system to make decisions based on high-level semantic understanding, rather than relying solely on color information. This semantics-aware approach is more robust in distinguishing between static and transient objects than traditional color-based methods.